OpenStudy (anonymous):

Simplify sqrt((t-7)^3)/t-7

OpenStudy (anonymous):

=sqrt(t-7)

OpenStudy (anonymous):

all of the numerator is to the third power

OpenStudy (anonymous):

\[(t-7)^{3/2} \over (t-7)\]

OpenStudy (anonymous):

3/2 - 1 =1/2

OpenStudy (anonymous):

@purplec16 sqrt(t-7) is the simplified ans.

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

i get it thank you so much

OpenStudy (anonymous):

no probs :)

OpenStudy (anonymous):

wait...that cant be the answer cause its a limit and we dont want it to equal to 0

OpenStudy (anonymous):

wait...that cant be the answer cause its a limit and we dont want it to equal to 0

OpenStudy (amistre64):

limits are allowed to be zero

OpenStudy (anonymous):

step1: sqrt((t-7)^3)/t-7 step2: ((t-7)^(3/2))/t-7 step3: ((t-7)^(3/2))*((t-7)^-1) (because 1/a=a^-1) step4: (t-7)^((3/2)-1) (because (a^m)*(a^n)=a^(m+n)) step5: (t-7)^(1/2) step5: sqrt(t-7)

OpenStudy (radar):

Problem states to simplify (t-7)^1/2

OpenStudy (anonymous):

but we want to solve it so that it is not 0

OpenStudy (anonymous):

but we want to solve it so that it is not 0

OpenStudy (amistre64):

no, we want to manipulate it so that we have no zero in a denominator

OpenStudy (amistre64):

since the bottom cancels with the top; all thats left is a simple ordinary 0

OpenStudy (anonymous):

hmm i as just going to say that

OpenStudy (anonymous):

was*